Identify as an increase or decrease. Then find the percent of increase or decrease. If necessary, round to the nearest tenth of a percent.
Original: 200
New: 170

Answer by math_tutor2020(3816)


Answer: 15% decrease


Work Shown:
a = old value = 200
b = new value = 170
c = change in values
c = b-a
c = 170-200
c = -30
The negative c value tells us that we have a percent decrease.

Then divide over the starting value a = 200
c/a = -30/200 = -0.15
Move the decimal point two spots to the right to go from 0.15 to 15%
Or multiply 0.15 with 100% to get 0.15*100% = 15%

We have a decrease of 15%

-------------------------------

Another approach:
a = old value = 200
b = new value = 170
b/a = ratio of new over old
b/a = 170/200 = 0.85

Subtract off 1 to get 0.85-1 = -0.15 which then converts to -15%
The negative percentage indicates percent decrease.

Here's a handy calculator to check your work.
https://www.calculatorsoup.com/calculators/algebra/percent-change-calculator.php